OpenStudy (anonymous):

How do i determine the mass of a planet when the radius is 3(10^7) and the gravitational field strength is 8g(N kg^-1)

OpenStudy (mos1635):

g1=G*M1/r^2 for planet go=G*Me/Re^2 for earth divide .....supstitude g1=8*go....solve to M1...

OpenStudy (anonymous):

pls can we get a clear solution to this problem @experimentX @shivam_bhalla

OpenStudy (experimentx):

M1 = Mass of the planet r = Radius of the planet Me = Mass of the Earth Re = Radius of the Earth g1 = Gravitational field strength of the planet go = Gravitational field strength of the earth These are the variables in the above equation.

OpenStudy (anonymous):

G is suppose to be Universal G constant then M1 can be found..then why do we need g1=8*go

OpenStudy (experimentx):

makes things much more faster!!!

OpenStudy (anonymous):

u mean g1 for any planet is 8times that of earth?

OpenStudy (experimentx):

yeah ... that what the question says. and you know the radius of earth.